# The Unmet Promise of Synthetic Training Images: Using Retrieved Real Images Performs Better

### v2

@@ Abstract
Generative text-to-image models enable us to synthesize unlimited amounts of images in a controllable manner, spurring many recent efforts to train vision models with synthetic data. However, every synthetic image ultimately originates from the upstream data used to train the generator. What additional value does the intermediate generator provide over directly training on relevant parts of the upstream data? Grounding this question in the setting of image classification, we compare finetuning on task-relevant, targeted synthetic data generated by Stable Diffusion—a generative model trained on the LAION-2B dataset—against finetuning on targeted real images retrieved directly from LAION-2B. We show that while synthetic data can benefit some downstream tasks, it is universally matched or outperformed by real data from the simple retrieval baseline. Our analysis suggests that this underperformance is partially due to generator artifacts and inaccurate task-relevant visual details in the synthetic images. Overall, we argue that retrieval is a critical baseline to consider when training with synthetic data—a baseline that current methods do not yet surpass. We release code, data, and models at https://github.com/scottgeng00/unmet-promise.
@@ 1 Introduction
Analogously, in computer vision—the focus of our research—many recent works train models on synthetic images from modern text-to-image generators, aiming to achieve state-of-the-art visual recognition performance  [54, 61, 3, 23, 24]. For example, SynCLR [60] cleverly prompts Stable Diffusion for synthetic images tailored to pre-specified downstream image recognition domains; they find that a CLIP-like model trained from scratch on the resulting targeted synthetic images can outperform CLIP trained on LAION-2B, a significantly larger untargeted dataset of real images. This result is quite surprising. Stable Diffusion is also trained on LAION-2B, so by the data processing inequality, the synthetic images it generates cannot contain any additional information over the images in LAION-2B. Yet, training on these derivative synthetic images appears to outperform training directly on LAION-2B. How do we make sense of these additional gains?
In this paper, we argue that the performance gained by training on generated synthetic images needs to be contextualized against a critical baseline often missing in prior work: training on real images from the generative model’s pretraining data. In particular, prior work has often compared task-targeted synthetic images to general, untargeted real images, thereby entangling the effects of training on synthetic versus real images with the effects of targeted versus general data collection. However, these variables are not intrinsically conflated. Any generative model we use to synthesize images fundamentally derives from its upstream training data. Instead of using that upstream data to train an intermediate generative model and synthesize targeted synthetic images, we can alternatively seek to directly retrieve targeted real images from the upstream source (Figure 1). By comparing synthetic training data against this retrieval baseline, we isolate the value added by the generative model.
@@
Overall, we conclude that retrieval is a critical baseline to consider when evaluating the true added utility of generated synthetic training data. Our goal is not to make normative claims about whether synthetic data will ever surpass this standard, but to contribute a simple baseline to aim for, and a clean set of experiments to explicitly measure progress towards surpassing it. For instance, by conceptualizing retrieval from a generator’s training data as a strong alternative to synthesizing data, a natural future direction for improving synthetic training data is to synthesize image compositions that are explicitly absent from the generator’s upstream training set. Images generated in this manner may offer unique value beyond what can be retrieved from the training data. Finally, in settings where the upstream dataset of a generator is unavailable altogether (e.g., due to privacy concerns, due to proprietary data, or due to download bandwidth restrictions), the retrieval baseline is unrealizable by assumption; synthetic data therefore retains strong utility for distilling knowledge from generative models and for privacy preservation. We release all code, models, and over 1TB of generated images to guide future work (https://github.com/scottgeng00/unmet-promise).
@@ 2 Related Work
Synthetic data has been widely explored in the context of many machine learning problems [66, 29, 25, 32, 59, 56, 5, 39]. For example, in NLP, synthetic data generated from strong large language models [46, 11] has been used to distill instruction-following behavior [66] and task-specific knowledge [30] into smaller models. In computer vision, prior works have sought to use synthetic data to improve the state-of-the-art across a breadth of visual tasks, such as object detection [47, 31], semantic segmentation [53, 51, 8], and optical flow [15]. Traditionally, this synthetic training data has been sourced from expert-crafted simulation and rendering pipelines [48, 15, 51, 8, 53, 31]. Recent advances in text-to-image synthesis via diffusion models [57, 27, 52] are changing this paradigm, inspiring a new wave of works that seek to train visual models on synthetic data algorithmically sampled from modern generative models [24, 61, 54, 23, 72]. This structural shift in the source of synthetic images from expert-supervised programmatic simulation to a learned generator that itself derives supervision from upstream data raises a critical question: does the intermediate step of training a generator and sampling synthetic data provide any gains over simply training on the upstream data directly? Our work formalizes and empirically grounds this question, contributing experiments and baselines to rigorously measure the benefits of training on modern data-derived synthetic data.
@@
Recent studies [72, 7] similarly use retrieval from generative model training data as a baseline for synthetic training images. For example, [72] adapts Stable Diffusion with real medical images and then generates synthetic medical images to train a ResNet from scratch; the resulting ResNet outperforms a ResNet trained on LAION-retrieved medical images. Consistent with our work, [72, 7] find that in the low-data regime of general image recognition tasks (e.g., subsampled ImageNet-1K, STL-10), augmenting real datasets with LAION-retrieved images outperforms augmenting with Stable Diffusion generated images when training ResNets from scratch. [7] further finds that this performance gap in the low-data regime persists across many synthetic image generation methods, including when they finetune Stable Diffusion with task-specific real data. In the medium-data regime (e.g. full ImageNet-1k), [7] finds augmenting with synthetic data matches but does not outperform augmenting with retrieved data. Our work generalizes the shared motivation of measuring the utility of synthetic training images to the modern data-rich regime and a wide range of visual classification tasks. We center our experiments on the scaling behavior of synthetic versus retrieved data for finetuning a web-scale pretrained CLIP model on general image recognition (ImageNet-1K), fine-grained recognition (FGVC-Aircraft, Flowers102, StanfordCars), and abstract concepts (DTD). We show that retrieved data often strictly outperforms synthetic data even at large data scales. We further contribute analyses to unpack the consistently observed performance gap between synthetic and retrieved real training images.
@@ 3.1 Sourcing data by generating synthetic images
We follow SynCLR [60], a representative method for curating synthetic training data from off-the-shelf text-to-image models. Given the set of visual class names ${\mathcal{C}}$, we first synthesize a large corpus of corresponding image captions by prompting a large language model (details in Appendix B.1). For example, if the class name $c\in{\mathcal{C}}$ is “rose,” then a generated caption might be “a close-up of a pink rose in bloom.” We then use these captions as input for a text-to-image generator $G$ trained on the upstream data ${\mathcal{D}}$, yielding a large set of synthesized images $\widetilde{x_{i}}$. Each image is assigned a class label $y_{i}$ based on the class name $c$ used to synthesize its caption. These synthetic images and labels $\{(\widetilde{x_{i}},y_{i})\}$ form our curated dataset ${\mathcal{D}_{\mathcal{C}}^{(\text{synthetic})}}$.
@@ 3.3 Additional data filtering and postprocessing
Data filtering has been shown to improve training performance for both real and synthetic data and is widely used in practice [20, 24]. We filter both our synthetic and retrieved datasets. Given a curated dataset ${\mathcal{D}_{\mathcal{C}}}$, we compute the CLIP similarity of each $x_{i}\in{\mathcal{D}_{\mathcal{C}}}$ with text corresponding to its assigned label $y_{i}$ (e.g., “a photo of {class name}”), constructed using the CLIP zero-shot classification templates [49]. When there are multiple templates for a given class, we aggregate by taking the maximum similarity across templates. We keep the top 30% of images by aggregate similarity. Intuitively, filtering helps remove generated and retrieved images with class-misaligned content. For example, an image labeled “dog” but without any dogs present (i.e., due to retrieval or generation errors) would receive a lower CLIP similarity score and thus likely be discarded.
@@ 4 Main Experiments
Benchmarks. We focus on five downstream tasks where synthetic data has shown promise compared to similar scale untargeted real data [60]. We select (a) ImageNet-1K [14] and Describable Textures (DTD) [12] to evaluate recognition performance on broad categories and (b) FGVC-Aircraft [40], StanfordCars [35], and Oxford Flowers102 [45] to evaluate performance in fine-grained settings. We use standard pre-defined train, test, and validation splits when available, and otherwise randomly subset the training set to create missing train-validation splits (details in Appendix C.2).
@@
Finetuning data curation. For each downstream benchmark, we first curate an adaptation dataset ${\mathcal{D}_{\mathcal{C}}}$ (Section 3) by either (1) generating synthetic images with Stable Diffusion 1.5 [52], trained on the LAION-2B dataset [55], or (2) retrieving real images directly from LAION-2B. We treat the choice between our substring-based and semantic retrieval strategies as a hyperparameter, using downstream validation set accuracy to determine the best choice for each benchmark. Hyperparameters for retrieval are detailed in Appendix C.1.
Model adaptation and evaluation. We adapt a LAION-2B pretrained CLIP ViT-B/16 [10] image encoder by finetuning on the curated adaptation dataset ${\mathcal{D}_{\mathcal{C}}}$ with a cross-entropy classification loss for a pre-set number of epochs. To elucidate the scaling trends of synthetic and retrieved data, we finetune across an order of magnitude of different adaptation dataset scales, subsampled from the full targeted adaptation dataset ${\mathcal{D}_{\mathcal{C}}}$. We report zero-shot (ZS) and linear probing (LP) test set accuracy, using the benchmark train set to train the linear head. For both LP and ZS evaluation, we use the validation set to identify the best epoch and finetuning hyperparameters. For each data scale, we aggregate accuracy across the results of at least three random seeds, and report the standard deviation due to seed randomness. Additional training and full hyperparameter details are provided in Appendix C.3.
@@ 5 Analysis
In this section, we explore two questions to better understand our main results. First, what factors drive the underperformance of synthetic data? Second, do our findings hold under variations of our experimental setup? We focus our analysis experiments on ImageNet, to understand general image recognition performance, and FGVC-Aircraft, the sole benchmark where synthetic data exhibited strong positive log-linear scaling.
@@ 5.1 Why does synthetic data lag retrieved real data?
To disentangle the effect of low-level generator artifacts and visual content differences between synthetic and retrieved real images on downstream model performance, we trained on “hybrid” images that have similar semantic visual content as our retrieved real images but contain generator artifacts like our synthetic images. Following SDEdit [42], we use Stable Diffusion to synthetically perturb our retrieved images to introduce model-specific artifacts present in the synthetic images Stable Diffusion generates. Given a noise strength parameter $\gamma\in[0,1]$ and a retrieved image $x_{0}$, SDEdit adds Gaussian noise to $x_{0}$ according to timestep $t=\gamma$ of Stable Diffusion’s time-dependent forward process. We then denoise the noisy image using the same reverse diffusion process as in text-to-image generation, yielding a perturbed image $x^{(\gamma)}$ that looks semantically like $x_{0}$ while also containing Stable Diffusion-specific artifacts. Increasing $\gamma$ increases the amount of Gaussian noise added to $x_{0}$, thereby increasing the severity of visual artifacts introduced in the resulting $x^{(\gamma)}$ (see Appendix D.1 for further details). In pseudocode,
@@
Figure 5: We finetune a pretrained CLIP model (dashed purple line) on retrieved real images that are synthetically perturbed (green circles) with Stable Diffusion to introduce generator artifacts. The perturbation strength is controlled by a parameter $\gamma\in[0,1]$ where larger $\gamma$ corresponds to stronger artifacts; within the gray-shaded region, the artifacts are strong enough to damage class-relevant details. Our results suggest that generator artifacts do contribute to synthetic data’s underperformance—with any amount of artifacts introduced, performance drops below training on retrieved images (dashed blue line). Moreover, differences in visual content between synthetic and retrieved images also matter; even with a relatively strong perturbation strength of $\gamma=0.5$, training on artifact-afflicted perturbed images that retain the semantic content of retrieved images outperforms training on synthetic images (dashed orange line).
@@ 6 Discussion
Our work sought to answer a key question: given that all model-generated synthetic images derive from the generator’s upstream training data, does training on synthetic images provide value over training on the parts of the upstream real data directly? We contribute a set of rigorous experiments to ground this question empirically, and discover that training on upstream real images collected via our simple retrieval baseline significantly outperforms training on synthetic images. Our initial question is answered negatively. We therefore argue that retrieval is a critical baseline to surpass in order to show value from synthetic training data, and encourage comparison against it in future research.
Importantly, we do not seek to make normative claims about whether training with synthetic images will ever surpass this baseline—future work may unlock gains that we have not yet found. As a first step, we contribute analyses of why synthetic training images underperform upstream real images, finding that both generator artifacts and semantic errors within synthetic images are key areas for future improvement. Furthermore, given that image retrieval is a strong alternative to image synthesis, a natural next step is to generate image compositions that are explicitly absent from the generator’s upstream training dataset; synthesizing these “missing" images may offer unique value beyond the existing upstream real images. Such an approach leverages the compositional generalization abilities of the generator, which recent research promisingly suggests may be stronger than the compositionality of a discriminative model trained on the same upstream data [37, 13].
@@
We graciously thank (in alphabetical order) Eric Frankel, Jacqueline He, Athena Tsu, and Rui Xin for their helpful comments and feedback. SG is supported by the NSF GRFP.

### v5

@@ Abstract
Generative text-to-image models enable us to synthesize unlimited amounts of images in a controllable manner, spurring many recent efforts to train vision models with synthetic data. However, every synthetic image ultimately originates from the upstream data used to train the generator. Does the intermediate generator provide additional information over directly training on relevant parts of the upstream data? Grounding this question in the setting of image classification, we compare finetuning on task-relevant, targeted synthetic data generated by Stable Diffusion—a generative model trained on the LAION-2B dataset—against finetuning on targeted real images retrieved directly from LAION-2B. We show that while synthetic data can benefit some downstream tasks, it is universally matched or outperformed by real data from the simple retrieval baseline. Our analysis suggests that this underperformance is partially due to generator artifacts and inaccurate task-relevant visual details in the synthetic images. Overall, we argue that targeted retrieval is a critical baseline to consider when training with synthetic data—a baseline that current methods do not yet surpass. We release code, data, and models at https://github.com/scottgeng00/unmet-promise.
@@ 1 Introduction
Analogously, in computer vision—the focus of our research—many recent works train models on synthetic images from modern text-to-image generators, aiming to achieve state-of-the-art visual recognition performance  [56, 63, 3, 23, 24]. For example, SynCLR [62] cleverly prompts Stable Diffusion for synthetic images tailored to manually-specified downstream image recognition domains; they find that a CLIP-like model trained from scratch on the resulting targeted synthetic images can outperform CLIP trained on LAION-2B, a significantly larger untargeted dataset of real images. This result is quite surprising. Stable Diffusion is also trained on LAION-2B, so by the data processing inequality, the synthetic images it generates cannot contain any additional information about LAION-2B over the original real LAION-2B images. Yet, training on these derivative synthetic images appears to outperform training directly on LAION-2B. How do we make sense of these additional gains? Does the generative model truly add useful information on top of its pretraining data?
In this paper, we argue that the performance gained from training on generated synthetic images needs to be contextualized against a critical baseline often missing in prior work: training on real images subselected from the generative model’s pretraining data. In particular, prior work has often compared task-targeted synthetic images to general, untargeted real images (e.g., full LAION-2B), thereby entangling the effects of training on synthetic versus real images with the effects of targeted versus general data collection. However, these variables are not intrinsically conflated. Any generative model we use to synthesize images fundamentally derives from its upstream training data. Instead of using that upstream data to train an intermediate generative model and synthesize targeted synthetic images, we can alternatively seek to directly identify targeted real images from the upstream source through retrieval (Figure 1). By comparing synthetic training data against this retrieval baseline, we isolate the value added by the generative model.
@@
Overall, we conclude that retrieval is a critical baseline to consider when evaluating the true added utility of generated synthetic training data. Our goal is not to make normative claims about whether synthetic data will ever surpass this standard, but to contribute a simple baseline to aim for, and a clean set of experiments to explicitly measure progress towards surpassing it. We are optimistic that synthetic data holds future promise; for instance, by conceptualizing retrieval from a generator’s training data as a strong alternative to synthesizing data, a natural future direction for improving synthetic training data is to synthesize image compositions that are explicitly absent from the generator’s upstream training set. Images generated in this manner may offer unique value beyond what can be retrieved from the training data. Finally, in settings where the upstream dataset of a generator is unavailable altogether (e.g., due to privacy concerns, due to proprietary data, or due to download bandwidth restrictions), the retrieval baseline is unrealizable by assumption; synthetic data therefore retains strong utility for distilling knowledge from generative models and for privacy preservation. We release all code, models, and over 1TB of generated images to guide future work (https://github.com/scottgeng00/unmet-promise).
@@ 2 Related Work
Synthetic data has been widely explored in the context of many machine learning problems [68, 29, 25, 32, 61, 58, 5, 40]. For example, in NLP, synthetic data generated from strong large language models [47, 11] has been used to distill instruction-following behavior [68] and task-specific knowledge [30] into smaller models. In computer vision, prior works have sought to use synthetic data to improve the state-of-the-art across a breadth of visual tasks, such as object detection [48, 31], semantic segmentation [54, 52, 8], and optical flow [15]. Traditionally, this synthetic training data has been sourced from expert-crafted simulation and rendering pipelines [49, 15, 52, 8, 54, 31]. Recent advances in text-to-image synthesis via diffusion models [59, 27, 53] are changing this paradigm, inspiring a new wave of works that seek to train visual models on synthetic data algorithmically sampled from large-scale generative models [24, 63, 56, 23, 74, 37]. Recent works have also sought to algorithmically sample synthetic data from diffusion models trained on data-scarce domains [55]. This structural shift in the source of synthetic images from expert-supervised programmatic simulation to a learned generator that itself derives supervision from upstream data raises a critical question: does the intermediate step of training a generator and sampling synthetic data provide any added useful information over simply training on relevant parts of the upstream data directly? Our work formalizes and empirically grounds this question, contributing experiments and baselines to rigorously measure the benefits of training on modern data-derived synthetic data.
@@
Recent studies [74, 7] similarly use retrieval from generative model training data as a baseline for synthetic training images. For example, [74] adapts Stable Diffusion with real medical images and then generates synthetic medical images to train a ResNet from scratch; the resulting ResNet outperforms a ResNet trained on LAION-retrieved medical images. Consistent with our work, [74, 7] find that in the low-data regime of general image recognition tasks (e.g., subsampled ImageNet-1K, STL-10), augmenting real datasets with LAION-retrieved images outperforms augmenting with Stable Diffusion generated images when training ResNets from scratch. [7] further finds that this performance gap in the low-data regime persists across many synthetic image generation methods, including when they finetune Stable Diffusion with task-specific real data. In the medium-data regime (e.g. full ImageNet-1k), [7] finds augmenting with synthetic data matches but does not outperform augmenting with retrieved data. Our work generalizes the shared motivation of measuring the utility of synthetic training images to the modern data-rich regime and a wide range of visual classification tasks.
@@ 3.1 Sourcing data by generating synthetic images
We follow SynCLR [62], a representative method for curating synthetic training data from off-the-shelf text-to-image models. Given the set of visual class names ${\mathcal{C}}$, we first synthesize a large corpus of corresponding image captions by prompting a large language model (details in Appendix C.1). For example, if the class name $c\in{\mathcal{C}}$ is “rose,” then a generated caption might be “a close-up of a pink rose in bloom.” We then use these captions as input for a text-to-image generator $G$ trained on the upstream data ${\mathcal{D}}$, yielding a large set of synthesized images $\widetilde{x_{i}}$. Each image is assigned a class label $y_{i}$ based on the class name $c$ used to synthesize its caption. These synthetic images and labels $\{(\widetilde{x_{i}},y_{i})\}$ form our curated dataset ${\mathcal{D}_{\mathcal{C}}^{(\text{synthetic})}}$.
@@ 3.3 Additional data filtering and postprocessing
Data filtering has been shown to improve training performance for both real and synthetic data and is widely used [20, 24]. We filter both our synthetic and retrieved datasets following current best filtering practices for synthetic image data [24]. Given a curated dataset ${\mathcal{D}_{\mathcal{C}}}$, we compute the CLIP similarity of each $x_{i}\in{\mathcal{D}_{\mathcal{C}}}$ with text corresponding to its assigned label $y_{i}$ (e.g., “a photo of {class name}”), constructed using the CLIP zero-shot classification templates [50]. When there are multiple templates for a given class, we aggregate by taking the maximum similarity across templates. We keep the top 30% of images per class by aggregate similarity. Intuitively, filtering helps remove generated and retrieved images with class-misaligned content. For example, an image labeled “dog” but without any dogs present (i.e., due to retrieval or generation errors) would receive a lower CLIP similarity score and likely be discarded. See Appendix C.2 for further discussion.
@@ 4 Main Experiments
Benchmarks. We focus on five downstream tasks where synthetic data has shown promise compared to similar scale untargeted real data [62]. We select (a) ImageNet-1K [14] and Describable Textures (DTD) [12] to evaluate recognition performance on broad categories and (b) FGVC-Aircraft [41], StanfordCars [35], and Oxford Flowers102 [46] to evaluate performance in fine-grained settings. We use standard pre-defined train, test, and validation splits when available, and otherwise randomly subset the training set to create missing train-validation splits (details in Appendix D.2).
@@
Finetuning data curation. For each downstream benchmark, we first curate an adaptation dataset ${\mathcal{D}_{\mathcal{C}}}$ (Section 3) by either (1) generating synthetic images with Stable Diffusion 1.5 [53], trained on the LAION-2B dataset [57], or (2) retrieving real images directly from LAION-2B. We treat the choice between our substring-based and semantic retrieval strategies as a hyperparameter, using downstream validation set accuracy to determine the best choice for each benchmark. Hyperparameters for retrieval are detailed in Appendix D.1.
Model adaptation and evaluation. We adapt a LAION-2B pretrained CLIP ViT-B/16 [10] image encoder by finetuning on the curated adaptation dataset ${\mathcal{D}_{\mathcal{C}}}$ with a cross-entropy classification loss for a pre-set number of epochs. To elucidate the scaling trends of synthetic and retrieved data, we finetune across an order of magnitude of different adaptation dataset scales, subsampled from the full targeted adaptation dataset ${\mathcal{D}_{\mathcal{C}}}$. We report zero-shot (ZS) and linear probing (LP) test set accuracy, using the benchmark train set to train the LP head. For both LP and ZS evaluation, we use the validation set to identify the best epoch and finetuning hyperparameters. For each data scale, we aggregate accuracy across the results of at least three random seeds, and report the standard deviation due to seed randomness. Please refer to Appendix D.3 for further evaluation details, and Appendix D.4 for further training and hyperparameter details.
@@ 5 Analysis
In this section, we explore two questions to better understand our main results. First, what factors drive the underperformance of synthetic data? Second, do our findings hold under variations of our experimental setup? We focus our analysis experiments on ImageNet, to understand general image recognition performance, and FGVC-Aircraft, the sole benchmark where synthetic data exhibited strong positive log-linear scaling. Additional analysis experiments are presented in Appendix B.
@@ 5.1 Why does synthetic data lag retrieved real data?
To disentangle the effect of low-level generator artifacts and visual content differences between synthetic and retrieved real images on downstream model performance, we trained on “hybrid” images that have similar semantic visual content as our retrieved real images but contain generator artifacts like our synthetic images. Following SDEdit [43], we use Stable Diffusion to synthetically perturb our retrieved images to introduce model-specific artifacts present in the synthetic images Stable Diffusion generates. Given a noise strength parameter $\gamma\in[0,1]$ and a retrieved image $x_{0}$, SDEdit adds Gaussian noise to $x_{0}$ according to timestep $t=\gamma$ of Stable Diffusion’s time-dependent forward process. We then denoise the noisy image using the same reverse diffusion process as in text-to-image generation, yielding a perturbed image $x^{(\gamma)}$ that looks semantically like $x_{0}$ while also containing Stable Diffusion-specific artifacts. Increasing $\gamma$ increases the amount of Gaussian noise added to $x_{0}$, thereby increasing the severity of visual artifacts introduced in the resulting $x^{(\gamma)}$ (see Appendix E.1 for further details). In pseudocode,
@@
Figure 5: We finetune a pretrained CLIP model (dashed purple line) on retrieved real images that are synthetically perturbed (green circles) with Stable Diffusion to introduce generator artifacts. The perturbation strength is controlled by a parameter $\gamma\in[0,1]$ where larger $\gamma$ introduces stronger artifacts; within the gray-shaded region, the artifacts are strong enough to damage class-relevant details. Our results suggest that generator artifacts do contribute to synthetic data’s underperformance—any artifact level causes performance to drop below training on retrieved images (dashed blue line). Moreover, differences in visual content between synthetic and retrieved images also matter; even with relatively strong perturbations ($\gamma=0.5$), training on artifact-afflicted perturbed images that retain the semantic content of retrieved images outperforms training on synthetic images (dashed orange line).
@@ 6 Discussion
Our work sought to answer a key question: given that all model-generated synthetic images derive from the generator’s upstream training data, does training on synthetic images provide gains over training on relevant parts of the upstream real data directly? In other words, do generative models add useful information on top of their own training data? Previous studies in the literature fall short of addressing this question fully, as they often compare task-targeted synthetic data to non-targeted subsets of the generator’s real training data, thus failing to control for the critical effect of targeted data collection. We propose the retrieval baseline as a means for controlling for this overlooked effect, and in doing so enable a new set of rigorous experiments to empirically ground our research question. We discover that training on upstream real images collected via our simple retrieval baseline significantly outperforms training on synthetic images. Our initial question is thus answered negatively: with existing best methods, synthetic data does not contain significant added information. We therefore argue that retrieval is a critical baseline to surpass in order to show value from synthetic training data, and encourage comparison against it in future research.
Importantly, we do not seek to make normative claims about whether training with synthetic images will ever surpass this baseline—future work may unlock gains that we have not yet found. As a first step, we contribute analyses of why synthetic training images underperform upstream real images, finding that both generator artifacts and semantic errors within synthetic images are key areas for future improvement. Furthermore, given that image retrieval is a strong alternative to image synthesis, a natural next step is to generate image compositions that are explicitly rare or absent from the generator’s upstream training dataset; we are optimistic that synthesizing these “missing" images may offer unique value beyond what is present in the existing upstream real images. Such an approach leverages the compositional generalization abilities of the generator, which recent research promisingly suggests may be stronger than the compositionality of a discriminative model trained on the same upstream data [38, 13].
@@
We graciously thank (in alphabetical order) Eric Frankel, Jacqueline He, Athena Tsu, and Rui Xin for their helpful comments and feedback. SG is supported by the NSF GRFP. PWK is supported by the Singapore National Research Foundation and the National AI Group in the Singapore Ministry of Digital Development and Information under the AI Visiting Professorship Programme (award number AIVP-2024-001).